# Algorithms and Complexity of Hedge Cluster Deletion Problems

### v1

@@ 4.1 Towards a constant-factor approximation
Here we consider approximating Hedge Cluster Deletion up to a constant factor, by restricting to certain class of hedge graphs. In order to achieve a constant factor approximation, we require a relationship between the hedges and the triangles of the underlying graph. In particular, we consider hedge graphs in which every triangle of the underlying graph is covered by at most two hedges. We refer to such hedge graphs as bi-hedge graphs. By Lemma 3.1, observe that Hedge Cluster Deletion remains NP-complete on bi-hedge graphs.
@@
To avoid heavy notation, we refer to a hedge of $\mathcal{E}(H)$ both as $x$ or $E_{x}$, whenever there is no ambiguity. For a hedge $x$, we denote by $SK_{3}(x)$ (single-$K_{3}$) the set of triangles $K_{3}=a,b,c$ in $H$ such that exactly one of the three edges belongs to $E_{x}$, that is $\{a,b\}\in E_{x}$ and $\{a,c\},\{b,c\}\notin E_{x}$. Since every triangle contains at most two hedges, notice that both $\{a,c\}$ and $\{b,c\}$ belong to a single hedge $E_{y}\neq E_{x}$. Let $D(x)$ be the set of hedges that appear in $SK_{3}(x)$. We say that a hedge $y$ dominates a hedge $x$ if there is a sequence of hedges $\langle y=h_{1},\ldots,h_{p}=x\rangle$ with $p\geq 1$ such that $h_{i}\in D(h_{i+1})$, for all $1\leq i<p$. We denote by $R(x)$ all hedges that dominate $x$. Notice that $x\in R(x)$. Figure 2 illustrates the corresponding notions.
@@
Let $H=(V,\mathcal{E})$ be a bi-hedge graph. From $H$, we construct a graph $G$ and a set $\mathcal{L}$ for Multi-Vertex Cover as follows. For every $E_{x}\in\mathcal{E}$, there is a corresponding vertex $x$ in $V(G)$ and there is an edge between $x$ and $y$ if the corresponding hedges $E_{x}$ and $E_{y}$ form a $P_{3}$-hedge in $H$. Moreover, for every vertex $v\in V(G)$, we have $L(v)=R(E_{v})$ for the corresponding hedge $E_{v}$. Observe that the construction takes polynomial time because $V(G)$ is exactly $|\mathcal{E}|$ and every $L(v)$ can be computed in polynomial time by Lemma 4.11 and has size at most $|\mathcal{E}|$. Notice also that, by construction, the statements for the hedges of Lemma 4.10 hold for the corresponding vertices in $G$. Thus, property (P1) is satisfied for the vertex-lists of $G$. We claim that there is a solution $U$ for Hedge Cluster Deletion of size at most $k$ if and only if there is a solution $S$ for Multi-Vertex Cover in $G$ of size at most $k$.
Let $U$ be a solution for Hedge Cluster Deletion with $k$ hedges. For every $P_{3}$-hedge, there exists a hedge $x$ such that $x\in U$. We claim that $S=U$ is a solution for Multi-Vertex Cover in $G$. By construction, for every $P_{3}$-hedge in $H$ there is an edge in $G$. Hence, we have that $S$ is a vertex cover for $G$, since there are no edges in GSG-S. Moreover, by Lemma 4.10 (ii) we have that $R(E_{x})\subseteq U$ for any hedge $E_{x}$. In terms of the vertex-lists of $G$, this means that $L(S)=S$, because $x\in L(x)$ for any vertex of $G$. Therefore, $S$ is a vertex cover for $G$ with $L(S)=S$ and $|S|\leq k$.
For the opposite direction, assume that there is a vertex cover $S$ of size at most $k$ for $G$ such that $L(S)=S$ and the vertex-lists satisfy property (P1). We claim that $U=S$ is a solution for Hedge Cluster Deletion. Assume for contradiction that there is at least one $P_{3}$-hedge in $H\setminus U$. Suppose that the underlying graph of $H\setminus U$ contains a $P_{3}=a,b,c$ induced by the vertices $a,b,c$. Let $E_{x}$ be a hedge of $H\setminus U$ such that $\{a,b\}\in E_{x}$. If $a,b,c$ induce a $P_{3}$ in $H$, then we have $\{b,c\}\in E_{y}$ with $E_{x}\neq E_{y}$, since there are no hedges in $H$ that span an internal $P_{3}$. Thus $x$ or $y$ belongs to $S$, because $S$ is a vertex cover of $G$. Hence, the vertices $a,b,c$ form a $K_{3}$ in $H$. As $H$ is bi-hedge, there are two hedges involved in the $K_{3}$. Since $\{a,c\}$ is not an edge of $H\setminus U$, the single hedge of the $K_{3}$ contains the edge $\{a,c\}$, so that $\{a,b,c\}\in SK(y)$. We conclude that $\{a,b\},\{b,c\}\in E_{x}$, $\{a,c\}\in E_{y}$, and $E_{y}\in U$. Thus, by definition we have $x\in D(y)$, which implies $x\in R(y)$. This leads to a contradiction by Lemma 4.10 (ii), since $y\in U$ and $R(y)\nsubseteq U$. Therefore, $U$ is a solution for Hedge Cluster Deletion. ∎
@@
Let $z$ be a vertex such that $z\in L(x)\cap L(y)$. Assume for contradiction that $z\notin S$. Since $\{x,y\}$ an edge of $G$, either $x$ or $y$ belongs to $S$. Without loss of generality, assume that $x\in S$. By the fact that $L(S)=S$ and $x\in S$, we have $L(x)\subseteq S$. The latter leads to a contradiction since $z\notin S$ and $z\in L(x)$. Hence, $L(x)\cap L(y)\subseteq S$. ∎
@@
Next, we show the minimality for $S^{\prime}$. Let $x\in S\cap V(G^{\prime})$. We show that there exists a vertex $w$ in $G^{\prime}$ such that $w\notin S^{\prime}$ with $\{x,w\}\in E(G^{\prime})$. By construction, $x\in S$. Since $S$ is minimal, we have two cases to consider. Assume that there is a vertex $y\notin S$ such that $\{x,y\}\in E(G)$. Then $y$ is a vertex of $G^{\prime}$, because for all vertices of $V(G)\setminus V(G^{\prime})$ we know that they belong to $S$ by 4.13. Thus $\{x,y\}\in E(G^{\prime})$ by construction, which means that $S^{\prime}$ is minimal because $y\notin S^{\prime}$ and $x\in S$. Assume that there is a vertex $y\in S$ such that $x\in L(y)$ and $\{y,z\}\in E(G)$ with $z\notin S$. Again, $z\notin S$ implies that $z\in V(G^{\prime})$. Now if $y\notin V(G^{\prime})$ then by property (P1) we have $x\notin V(G^{\prime})$, which is not possible. Thus $x,y,z$ are vertices of $G^{\prime}$, which implies that $\{x,z\}\in E(G^{\prime})$ by 4.15, since $x\in L(y)$. Therefore, $S^{\prime}$ is a minimal vertex cover for $G^{\prime}$.
For the opposite direction, assume a minimal vertex cover $S^{\prime}$ for $G^{\prime}$. We show that $S=S^{\prime}\cup\{z\,|\,z\in L(x)\cap L(y)\,\text{with }\{x,y\}\in E(G)\}$ is a minimal solution of Multi-Vertex Cover. By 4.13, all vertices of $L(x)\cap L(y)$ with $\{x,y\}\in E(G)$ belong to $S$. Since $S^{\prime}$ is a minimal vertex cover for $G^{\prime}$, there exists a vertex $x^{\prime}\in S^{\prime}$ and a vertex $y^{\prime}\notin S^{\prime}$ for some edge $\{x^{\prime},y^{\prime}\}\in G^{\prime}$. By the discussion above, we know that there is an edge $\{x,y\}\in(E(G^{\prime})\cap E(G))$ such that $x^{\prime}\in L(x)$ and $y^{\prime}\in L(y)$. We show that $x\in S$. To see this, observe that there is the edge $\{x,y^{\prime}\}\in E(G^{\prime})$ by construction. Since $y^{\prime}\notin S^{\prime}$ and $S^{\prime}$ is a vertex cover, we conclude that $x\in S^{\prime}$ which implies that $x\in S$. We further show that $y\notin S^{\prime}$. If $y\in S^{\prime}$, there is an edge $\{y,z\}\in E(G^{\prime})$ such that $z\notin S^{\prime}$ because $S^{\prime}$ is minimal. Since $y^{\prime}\in L(y)$, we have $\{y^{\prime},z\}\in E(G^{\prime})$. However, $y^{\prime},z\notin S^{\prime}$ leads to a contradiction, since $S^{\prime}$ is a vertex cover for $G^{\prime}$. Hence, $y\notin S^{\prime}$ which means that $y\notin S$. To conclude, for any edge $\{x,y\}$ of $E(G)$, we have $x\in S$ and $y\notin S$, or $x\in S$ and there is a vertex $x^{\prime}\in S$ such that $x\in L(x^{\prime})$ and $\{x^{\prime},y\}\in E(G)$ with $y\notin S$. This is exactly the definition of the minimal solution and we conclude that $S^{\prime}$ is indeed a minimal solution for Multi-Vertex Cover.
@@ 5 Acyclic hedge intersection graph
In our algorithm we consider mainly the hedge intersection graph $\mathcal{F}$ of a hedge graph $H$. To avoid heavy notation, we refer to the elements of $V(\mathcal{F})$ and $E(\mathcal{F})$, as vertices (that correspond to the hedges $\mathcal{E}(H)$) and edges. For a vertex $x\in V(\mathcal{F})$, we refer to $E_{x}$ as the corresponding hedge of $\mathcal{E}(H)$. By definition, an edge $\{x,y\}$ of $\mathcal{F}$ is characterized as follows: there are edges $\{u,v\}\in E_{x}$ and $\{v,w\}\in E_{y}$ in $H$. This means that there is either a $P_{3}$ in $H$ or a $K_{3}$ whenever $\{u,w\}$ belongs to a hedge of $H$. In the latter, observe that every triangle of the underlying graph contains at most two hedges, because $\mathcal{F}$ is acyclic. Thus, $\{u,w\}\in(E_{x}\cup E_{y})$.
@@
  • $E_{P}\cap E_{\triangle}$: $E_{x}$ and $E_{y}$ span both a $P_{3}$ and a $K_{3}$ in $H$.
We now borrow the notation given in Section 4.1. For any edge $\{x,y\}$ of $E_{\triangle}$ or $E_{P}\cap E_{\triangle}$, we know that $y\in D(x)$ or $x\in D(y)$ because there is a triangle with edges only of $E_{x}\cup E_{y}$. A vertex $x$ of $\mathcal{F}$ is called mixed vertex if $x$ is incident to an edge $\{x,y\}$ of $E_{P}\cap E_{\triangle}$ such that $x\in D(y)$. In what follows, to avoid repeating ourselves we let $U$ be a solution for Hedge Cluster Deletion on a hedge graph $H$ that admits an acyclic hedge intersection graph $\mathcal{F}$.
@@
Let $Z$ be the set of hedges that span an internal $P_{3}$ in $H$ and let $X_{P\cap\triangle}$ be the set of all mixed vertices of $\mathcal{F}$. Then, $Z\cup X_{P\cap\triangle}\subseteq U$.
@@
By 2.4, we know that $Z\cup U$. We now show that any mixed vertex $x$ belongs to $U$. Let $\{x,y\}$ be the edge of $E_{P}\cap E_{\triangle}$ incident to $x$. Since there is a $P_{3}$ spanned by edges of $E_{x}\cup E_{y}$, we know that $x$ or $y$ or both belong to $U$. Assume that $y\in U$. Then, we show that $x\in U$. To see this, observe that $x\in D(y)$ implies that there is a $K_{3}$ with two of its edges in $E_{x}$ and the other edge in $E_{y}$. Since $y\in U$ and $x\notin U$, we have an internal $P_{3}$ spanned by edges only of $E_{x}$. Therefore we conclude that $x\in U$. ∎
According to Lemma 5.1, we can safely remove the vertices of $Z\cup X_{P\cap\triangle}$ and consider the remaining graph. Let ${F}_{1}=\mathcal{F}-(Z\cup X_{P\cap\triangle})$. As ${F}_{1}$ is an induced subgraph of $\mathcal{F}$, we know that ${F}_{1}$ remains acyclic. Moreover observe that the edges of ${F}_{1}$ are partitioned into the sets $E_{P}$ and $E_{\triangle}$. Since $Z\cup X_{P\cap\triangle}$ are hedges of $H$, $H^{\prime}=H\setminus(Z\cup X_{P\cap\triangle})$ is a hedge-subgraph of $H$.
@@
Next suppose that $P_{3}$ is induced in $H^{\prime}\setminus U^{\prime}$, but not induced in $H^{\prime}$. This particularly means that $\{a,c\}$ being an edge in $H^{\prime}$. Since $E_{x}\notin U^{\prime}$, we have $\{a,c\}\in E_{y}$ with $E_{y}\in U^{\prime}$. If $\{b,c\}\notin E_{x}$ then $\{b,c\}\in E_{y}$ because there is a triangle in $H^{\prime}$ and $F_{1}$ is acyclic. This, however, contradicts the fact that $E_{y}\in U^{\prime}$. Thus $\{b,c\}\in E_{x}$. Now observe that $E_{x}$ and $E_{y}$ span a $K_{3}$ in $H$ which implies that $\{x,y\}$ is an edge of $E_{\triangle}$ or $E_{P}\cap E_{\triangle}$. If $\{x,y\}\in E_{P}\cap E_{\triangle}$ then $x\in D(y)$, so that $x\in X_{P\cap\triangle}$. This, however, implies that $x$ is not a vertex of $F_{2}$ due to Lemma 5.1. Thus $\{x,y\}\in E_{\triangle}$. In particular, $x$ dominates $y$ and both vertices belong to the same component of $F_{2}$ by Lemma 5.3. Since $S$ is a minimal vertex cover and $y\in S$, there is a vertex $z\notin S$ such that $\{y,z\}$ is an edge of $F_{2}$. Due to Lemma 5.4, there is an edge $\{x,z\}$ in $F_{2}$. Then, however we reach a contradiction to the fact that $S$ being a vertex cover of $F_{2}$ because $x,z\notin S$. Therefore, $U^{\prime}=S$ is a minimal solution for Hedge Cluster Deletion on $H^{\prime}$. ∎
@@
4:Partition the edges of $\mathcal{F}$ into the three sets: $_{P}, EE_{\triangle}$, $E_{P}\cap E_{\triangle}$
@@
7:Let $X_{P\cap\triangle}$ be the mixed vertices of $\mathcal{F}$
8:Compute the graph $H^{\prime}=H\setminus(Z\cup X_{P\cap\triangle})$ and update accordingly $R(x)$
9:Compute the components $V_{1},\ldots,V_{q}$ of $F_{1}=\mathcal{F}-(Z\cup X_{P\cap\triangle})$
@@
14:Return $U=Z\cup X_{P\cap\triangle}\cup U^{\prime}$
@@
We describe the steps in Algorithm 1. Given $H$, we begin by computing the hedge intersection graph $\mathcal{F}$ in $O(\ell m+n)$ time, where $\ell,m,n$ stand for the number of hedges, edges, and vertices, respectively. We also enumerate all $P_{3}$ and $K_{3}$ of the underlying graph in $O(n^{3})$ time. Having the enumerated subgraphs, we characterize each edge of $\mathcal{F}$ as $_{P}, EE_{\triangle}$, or $E_{P}\cap E_{\triangle}$. For the domination relation between the hedges, we can compute $R(x)$ in polynomial time by Lemma 4.11.
We collect all hedges that span an internal $P_{3}$ and all mixed vertices that correspond to $Z\cup X_{P\cap\triangle}$. We compute the graph $F_{2}$ in $O(\ell^{2})$ time and then compute a minimum vertex cover $U^{\prime}$ of $F_{2}$ in polynomial time [matching80, flow13], since $F_{2}$ is a bipartite graph by Lemma 5.6. We finally return the set $U^{\prime}\cup Z\cup X_{P\cap\triangle}$. Correctness follows from Lemmas 5.1 and 5.5. Therefore, all steps can be carried out in polynomial time. ∎

### v2

@@ 5 Towards a constant-factor approximation
Here we consider approximating Hedge Cluster Deletion up to a constant factor, by restricting to a certain class of hedge graphs. In order to achieve a constant factor approximation, we require a relationship between the hedges and the triangles of the underlying graph. In particular, we consider hedge graphs in which every triangle of the underlying graph is covered by at most two hedges. We refer to such hedge graphs as bi-hedge graphs. By Lemma 3.1, observe that Hedge Cluster Deletion remains NP-complete on bi-hedge graphs.
@@
To avoid heavy notation and without affecting consistency, we refer to a hedge of $\mathcal{E}(H)$ both as $x$ or $E_{x}$, whenever there is no ambiguity. For a hedge $x$, we denote by $SK_{3}(x)$ (single-$K_{3}$) the set of triangles $K_{3}=(a,b,c)$ in $H$ such that exactly one of the three edges belongs to $E_{x}$, that is $\{a,b\}\in E_{x}$ and $\{a,c\},\{b,c\}\notin E_{x}$. Since every triangle contains at most two hedges, notice that both $\{a,c\}$ and $\{b,c\}$ belong to a single hedge $E_{y}\neq E_{x}$. Let $D(x)$ be the set of hedges that appear in $SK_{3}(x)$. We say that a hedge $y$ dominates a hedge $x$ if there is a sequence of hedges $\langle y=h_{1},\ldots,h_{p}=x\rangle$ with $p\geq 1$ such that $h_{i}\in D(h_{i+1})$, for all $1\leq i<p$. We denote by $R(x)$ all hedges that dominate $x$. Notice that $x\in R(x)$. Figure 3 illustrates the corresponding notions.
@@
Let $H=(V,\mathcal{E})$ be a bi-hedge graph. From $H$, we construct a graph $G$ and a set $\mathcal{L}$ for Multi-Vertex Cover as follows. For every $E_{x}\in\mathcal{E}$, there is a corresponding vertex $x$ in $V(G)$ and there is an edge between $x$ and $y$ if the corresponding hedges $E_{x}$ and $E_{y}$ form a $P_{3}$ in $H$. Moreover, for every vertex $v\in V(G)$, we have $L(v)=R(E_{v})$ for the corresponding hedge $E_{v}$. Observe that the construction takes polynomial time because $V(G)$ is exactly $|\mathcal{E}|$ and every $L(v)$ can be computed in polynomial time by Lemma 5.2 and has size at most $|\mathcal{E}|$. Notice also that, by construction, the statements for the hedges of Lemma 5.1 hold for the corresponding vertices in $G$. Thus, property (P1) is satisfied for the vertex-lists of $G$. We claim that there is a solution $U$ for Hedge Cluster Deletion of size at most $k$ if and only if there is a solution $S$ for Multi-Vertex Cover in $G$ of size at most $k$.
Let $U$ be a solution for Hedge Cluster Deletion with $k$ hedges. For every $P_{3}$, there exists a hedge $x$ such that $x\in U$. We claim that $S=U$ is a solution for Multi-Vertex Cover in $G$. By construction, for every $P_{3}$ in $H$ there is an edge in $G$. Hence, we have that $S$ is a vertex cover for $G$, since there are no edges in GSG-S. Moreover, by Lemma 5.1 (ii) we have that $R(E_{x})\subseteq U$ for any hedge $E_{x}$. In terms of the vertex-lists of $G$, this means that $L(S)=S$, because $x\in L(x)$ for any vertex of $G$. Therefore, $S$ is a vertex cover for $G$ with $L(S)=S$ and $|S|\leq k$.
For the opposite direction, assume that there is a vertex cover $S$ of size at most $k$ for $G$ such that $L(S)=S$ and the vertex-lists satisfy property (P1). We claim that $U=S$ is a solution for Hedge Cluster Deletion. Assume for contradiction that there is at least one $P_{3}$ in $H\setminus U$. Suppose that the underlying graph of $H\setminus U$ contains a $P_{3}=(a,b,c)$ induced by the vertices $a,b,c$. Let $E_{x}$ be a hedge of $H\setminus U$ such that $\{a,b\}\in E_{x}$. If $a,b,c$ induce a $P_{3}$ in $H$, then we have $\{b,c\}\in E_{y}$ with $E_{x}\neq E_{y}$, since there are no hedges in $H$ that span an internal $P_{3}$. Thus $x$ or $y$ belongs to $S$, because $S$ is a vertex cover of $G$. Hence, the vertices $a,b,c$ form a $K_{3}$ in $H$. As $H$ is bi-hedge, there are two hedges involved in the $K_{3}$. Since $\{a,c\}$ is not an edge of $H\setminus U$, the single hedge of the $K_{3}$ contains the edge $\{a,c\}$, so that $\{a,b,c\}\in SK(y)$. We conclude that $\{a,b\},\{b,c\}\in E_{x}$, $\{a,c\}\in E_{y}$, and $E_{y}\in U$. Thus, by definition we have $x\in D(y)$, which implies $x\in R(y)$. This leads to a contradiction by Lemma 5.1 (ii), since $y\in U$ and $R(y)\nsubseteq U$. Therefore, $U$ is a solution for Hedge Cluster Deletion. ∎
@@
Let $z$ be a vertex such that $z\in L(x)\cap L(y)$. Assume for contradiction that $z\notin S$. Since $\{x,y\}$ is an edge of $G$, either $x$ or $y$ belongs to $S$. Without loss of generality, assume that $x\in S$. By the fact that $L(S)=S$ and $x\in S$, we have $L(x)\subseteq S$. The latter leads to a contradiction since $z\notin S$ and $z\in L(x)$. Hence, $L(x)\cap L(y)\subseteq S$. ∎
@@
Next, we show the minimality for $S^{\prime}$. Let $x\in S\cap V(G^{\prime})$. We show that there exists a vertex $w$ in $G^{\prime}$ such that $w\notin S^{\prime}$ with $\{x,w\}\in E(G^{\prime})$. By construction, $x\in S$. Since $S$ is minimal, we have two cases to consider. Assume that there is a vertex $y\notin S$ such that $\{x,y\}\in E(G)$. Then $y$ is a vertex of $G^{\prime}$, because for all vertices of $V(G)\setminus V(G^{\prime})$ we know that they belong to $S$ by 5.4. Thus $\{x,y\}\in E(G^{\prime})$ by construction, which means that $S^{\prime}$ is minimal because $y\notin S^{\prime}$ and $x\in S$. Assume that there is a vertex $y\in S$ such that $x\in L(y)$ and $\{y,z\}\in E(G)$ with $z\notin S$. Again, $z\notin S$ implies that $z\in V(G^{\prime})$. Now if $y\notin V(G^{\prime})$ then by property (P1) we have $x\notin V(G^{\prime})$, which is not possible. Thus $x,y,z$ are vertices of $G^{\prime}$, which implies that $\{x,z\}\in E(G^{\prime})$ by 5.6, since $x\in L(y)$. Therefore, $S^{\prime}$ is a minimal vertex cover for $G^{\prime}$.
For the opposite direction, assume a minimal vertex cover $S^{\prime}$ for $G^{\prime}$. We show that $S=S^{\prime}\cup\{z\,|\,z\in L(x)\cap L(y)\,\text{with }\{x,y\}\in E(G)\}$ is a minimal solution of Multi-Vertex Cover. By 5.4, all vertices of $L(x)\cap L(y)$ with $\{x,y\}\in E(G)$ belong to $S$. Since $S^{\prime}$ is a minimal vertex cover for $G^{\prime}$, there exists a vertex $x^{\prime}\in S^{\prime}$ and a vertex $y^{\prime}\notin S^{\prime}$ for some edge $\{x^{\prime},y^{\prime}\}\in G^{\prime}$. By the discussion above, we know that there is an edge $\{x,y\}\in(E(G^{\prime})\cap E(G))$ such that $x^{\prime}\in L(x)$ and $y^{\prime}\in L(y)$. We show that $x\in S$. To see this, observe that there is the edge $\{x,y^{\prime}\}\in E(G^{\prime})$ by construction. Since $y^{\prime}\notin S^{\prime}$ and $S^{\prime}$ is a vertex cover, we conclude that $x\in S^{\prime}$ which implies that $x\in S$. We further show that $y\notin S^{\prime}$. If $y\in S^{\prime}$, there is an edge $\{y,z\}\in E(G^{\prime})$ such that $z\notin S^{\prime}$ because $S^{\prime}$ is minimal. Since $y^{\prime}\in L(y)$, we have $\{y^{\prime},z\}\in E(G^{\prime})$. However, $y^{\prime},z\notin S^{\prime}$ leads to a contradiction, since $S^{\prime}$ is a vertex cover for $G^{\prime}$. Hence, $y\notin S^{\prime}$ which means that $y\notin S$. To conclude, for any edge $\{x,y\}$ of $E(G)$, we have $x\in S$ and $y\notin S$, or $x\in S$ and there is a vertex $x^{\prime}\in S$ such that $x\in L(x^{\prime})$ and $\{x^{\prime},y\}\in E(G)$ with $y\notin S$. This is exactly the definition of the minimal solution and we conclude that $S^{\prime}$ is indeed a minimal solution for Multi-Vertex Cover.
@@ 6 Acyclic hedge intersection graph
In our algorithm we consider mainly the hedge intersection graph $\mathcal{F}$ of a hedge graph $H$. We refer to the elements of $V(\mathcal{F})$ and $E(\mathcal{F})$, as vertices (that correspond to the hedges $\mathcal{E}(H)$) and edges. For a vertex $x\in V(\mathcal{F})$, we refer to $E_{x}$ as the corresponding hedge of $\mathcal{E}(H)$. By definition, an edge $\{x,y\}$ of $\mathcal{F}$ is characterized as follows: there are edges $\{u,v\}\in E_{x}$ and $\{v,w\}\in E_{y}$ in $H$. This means that there is either a $P_{3}$ in $H$ or a $K_{3}$ whenever $\{u,w\}$ belongs to a hedge of $H$. In the latter, observe that every triangle of the underlying graph contains at most two hedges, because $\mathcal{F}$ is acyclic. Thus, $\{u,w\}\in(E_{x}\cup E_{y})$.
@@
  • $E_{P,\triangle}$: $E_{x}$ and $E_{y}$ span both a $P_{3}$ and a $K_{3}$ in $H$.
We now borrow the notation given in Section 5. For any edge $\{x,y\}$ of $E_{\triangle}$ or $E_{P,\triangle}$, we know that $y\in D(x)$ or $x\in D(y)$ because there is a triangle with edges only of $E_{x}\cup E_{y}$. A vertex $x$ of $\mathcal{F}$ is called mixed vertex if $x$ is incident to an edge $\{x,y\}$ of $E_{P,\triangle}$ such that $x\in D(y)$. In what follows, to avoid repeating ourselves we let $U$ be a solution for Hedge Cluster Deletion on a hedge graph $H$ that admits an acyclic hedge intersection graph $\mathcal{F}$.
@@
Let $Z$ be the set of hedges that span an internal $P_{3}$ in $H$ and let $X_{P,\triangle}$ be the set of all mixed vertices of $\mathcal{F}$. Then, $Z\cup X_{P,\triangle}\subseteq U$.
@@
By 2.4, we know that $Z\subseteq U$. We now show that any mixed vertex $x$ belongs to $U$. Let $\{x,y\}$ be the edge of $E_{P,\triangle}$ incident to $x$. Since there is a $P_{3}$ spanned by edges of $E_{x}\cup E_{y}$, we know that $x$ or $y$ or both belong to $U$. Assume that $y\in U$. Then, we show that $x\in U$. To see this, assume for contradiction that $x\notin U$. Observe that $x\in D(y)$ implies that there is a $K_{3}$ with two of its edges in $E_{x}$ and the other edge in $E_{y}$. Since $y\in U$ and $x\notin U$, we have an internal $P_{3}$ spanned by edges only of $E_{x}$. Therefore we conclude that $x\in U$. ∎
According to Lemma 6.1, we can safely remove the vertices of $Z\cup X_{P,\triangle}$ and consider the remaining graph. Let ${F}_{1}=\mathcal{F}-(Z\cup X_{P,\triangle})$. As ${F}_{1}$ is an induced subgraph of $\mathcal{F}$, we know that ${F}_{1}$ remains acyclic. Moreover observe that the edges of ${F}_{1}$ are partitioned into the sets $E_{P}$ and $E_{\triangle}$. Since $Z\cup X_{P,\triangle}$ are hedges of $H$, $H^{\prime}=H\setminus(Z\cup X_{P,\triangle})$ is a hedge-subgraph of $H$.
@@
Next suppose that $P_{3}$ is induced in $H^{\prime}\setminus U^{\prime}$, but not induced in $H^{\prime}$. This particularly means that $\{a,c\}$ being an edge in $H^{\prime}$. Since $E_{x}\notin U^{\prime}$, we have $\{a,c\}\in E_{y}$ with $E_{y}\in U^{\prime}$. If $\{b,c\}\notin E_{x}$ then $\{b,c\}\in E_{y}$ because there is a triangle in $H^{\prime}$ and $F_{1}$ is acyclic. This, however, contradicts the fact that $E_{y}\in U^{\prime}$. Thus $\{b,c\}\in E_{x}$. Now observe that $E_{x}$ and $E_{y}$ span a $K_{3}$ in $H$ which implies that $\{x,y\}$ is an edge of $E_{\triangle}$ or $E_{P,\triangle}$. If $\{x,y\}\in E_{P,\triangle}$ then $x\in D(y)$, so that $x\in X_{P,\triangle}$. This, however, implies that $x$ is not a vertex of $F_{2}$ due to Lemma 6.1. Thus $\{x,y\}\in E_{\triangle}$. In particular, $x$ dominates $y$ and both vertices belong to the same component of $F_{2}$ by Lemma 6.3. Since $S$ is a minimal vertex cover and $y\in S$, there is a vertex $z\notin S$ such that $\{y,z\}$ is an edge of $F_{2}$. Due to Lemma 6.4, there is an edge $\{x,z\}$ in $F_{2}$. Then, however we reach a contradiction to the fact that $S$ being a vertex cover of $F_{2}$ because $x,z\notin S$. Therefore, $U^{\prime}=S$ is a minimal solution for Hedge Cluster Deletion on $H^{\prime}$. ∎
@@
4:Partition the edges of $\mathcal{F}$ into the three sets: $_{P}, EE_{\triangle}$, $E_{P,\triangle}$
@@
7:Let $X_{P,\triangle}$ be the mixed vertices of $\mathcal{F}$
8:Compute the graph $H^{\prime}=H\setminus(Z\cup X_{P,\triangle})$ and update accordingly $R(x)$
9:Compute the components $V_{1},\ldots,V_{q}$ of $F_{1}=\mathcal{F}-(Z\cup X_{P,\triangle})$
@@
14:Return $U=Z\cup X_{P,\triangle}\cup U^{\prime}$
@@
We describe the steps in Algorithm 1. Given $H$, we begin by computing the hedge intersection graph $\mathcal{F}$ in $O(\ell m+n)$ time, where $\ell,m,n$ stand for the number of hedges, edges, and vertices, respectively. We also enumerate all $P_{3}$ and $K_{3}$ of the underlying graph in $O(n^{3})$ time. Having the enumerated subgraphs, we characterize each edge of $\mathcal{F}$ as $_{P}, EE_{\triangle}$, or $E_{P,\triangle}$. For the domination relation between the hedges, we can compute $R(x)$ in polynomial time by Lemma 5.2.
We collect all hedges that span an internal $P_{3}$ and all mixed vertices that correspond to $Z\cup X_{P,\triangle}$. We compute the graph $F_{2}$ in $O(\ell^{2})$ time and then compute a minimum vertex cover $U^{\prime}$ of $F_{2}$ in polynomial time [matching80, flow13], since $F_{2}$ is a bipartite graph by Lemma 6.6. We finally return the set $U^{\prime}\cup Z\cup X_{P,\triangle}$. Correctness follows from Lemmas 6.1 and 6.5. Therefore, all steps can be carried out in polynomial time. ∎